OpenStudy (anonymous):

x-2/3=5/6 solve using the addition principle dont forget to check! I have 3x-2/3=5/6, 6(3x-2/3)=6*5/6, 18x-4=5, 4+18-4=5+4, 18x=9 please help with next step

OpenStudy (phi):

do you start with \[\frac{(x-2)}{3}= \frac{5}{6} \]?

OpenStudy (phi):

whatever it is, if you multiply by 3, you must: 1) multiply both sides by 3 2) multiply every term by 3

OpenStudy (anonymous):

start with x-2/3=5/6

OpenStudy (anonymous):

Thanks for the help